# Continuous Diffusion for Mixed-Type Tabular Data

### v2

@@ 1 Introduction
Score-based generative models (Song et al., 2021), also termed diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020), have shown remarkable potential for the generation of images (Dhariwal & Nichol, 2021; Rombach et al., 2022), videos (Ho et al., 2022), text (Li et al., 2022; Dieleman et al., 2022; Wu et al., 2023), molecules (Hoogeboom et al., 2022), and many other highly complex data structures with continuous features. The framework has since been adapted to categorical data in various ways, including discrete diffusion processes (Austin et al., 2021; Hoogeboom et al., 2021), diffusion in continuous embedding space (Dieleman et al., 2022; Li et al., 2022; Regol & Coates, 2023; Strudel et al., 2022), and others (Campbell et al., 2022; Meng et al., 2022; Sun et al., 2023). Diffusion models which include both, continuous and categorical features alike, build directly on advances from the image domain (Kim et al., 2023; Kotelnikov et al., 2023; Lee et al., 2023; Jolicoeur-Martineau et al., 2024) and thus, are not designed to deal with challenges specific to mixed-type data: The different diffusion processes are not aligned, and do not scale to larger datasets and/or features with a greater number of categories.
@@
A crucial component in score-based generative models is the noise schedule (Kingma et al., 2022; Chen et al., 2022; Chen, 2023; Jabri et al., 2022; Wu et al., 2023). Typical noise schedules for image and text data are designed to focus model capacity on the noise levels most important to sample quality (Nichol & Dhariwal, 2021; Karras et al., 2022), while others attempt to learn the optimal noise schedule (Dieleman et al., 2022; Kingma et al., 2022). For mixed-type tabular data, existing approaches often combine distinct diffusion processes for the continuous and discrete features to derive a joint model (Kotelnikov et al., 2023; Lee et al., 2023). However, noise schedules are not directly transferable from one data modality to another and therefore, using specifications from image or text domain models is not optimal: First, the inherently different diffusion processes make it difficult to balance the noise schedules across features and feature types, and negatively affect the allocation of model capacity across timesteps. Further, and most importantly, the domain, nature and marginal distribution can vary significantly across features (Xu et al., 2019). For instance, any two continuous features may be subject to different levels of discretization or different bounds, even after applying common data pre-processing techniques; and any two categorical features may differ in the number of categories, or the degree of imbalance. This high heterogeneity warrants a rethinking of fundamental parts of the diffusion framework, including the noise schedule and the effective combination of diffusion processes for different data types.
@@
We counteract the high feature heterogeneity inherent to data of mixed-type with distinct feature- or type-specific adaptive noise schedules. The learnable noise schedules allow the model to directly take feature or type heterogeneity into account during both training and generation, and thus avoid the reliance on image or text-specific noise schedules designs. This ensures a better allocation of the model’s capacity across features, feature types and timesteps, and guarantees high quality samples of tabular data. CDTD outperforms state-of-the-art baseline models across a diverse set of sample quality metrics, and computation time for data sets with an arbitrary number of categories and data points. Our experiments show that CDTD captures feature correlations exceptionally well, and that explicitly allowing for data-type heterogeneity in the noise schedules benefits sample quality.
@@
  • We propose a joint continuous diffusion model for both continuous and categorical features such that all noise distributions can be Gaussian.
  • We balance model capacity across continuous and categorical features with a novel and effective loss calibration.
  • We derive a functional form to learn adaptive noise schedules and to allow for an efficient, exact evaluation and thus, greatly extend the idea of timewarping.
@@
  • We boost the quality of the generated samples with adaptive, feature- or type-specific noise schedules.
  • Our novel combination of score estimation methods allows for the application of state-of-the-art techniques, such as classifier-free guidance and accelerated sampling, to mixed-type tabular data for the first time.
Figure 1: CDTD framework. Adaptive noise schedules are trained to fit the (possibly aggregated) MSE and CE losses and transform the uniform timestep $t$ to a potentially feature-specific noise level to diffuse (“noisify”) the scalar values (for continuous features) or the embeddings (for categorical features). Associated sampling processes are highlighted in orange. The approximated score functions are concatenated and passed to a blackbox ODE solver for sample generation.
@@ 2.1 Continuous Features
We denote the $i$-th continuous feature as $x_{\text{cont}}^{(i)}\in{\mathbb{R}}$ and the stacked feature vector as ${\mathbf{x}}_{0}\equiv{\mathbf{x}}_{\text{cont}}\in{\mathbb{R}}^{K_{\text{cont}}}$. Further, let $\{{\mathbf{x}}_{t}\}_{t=0}^{t=1}$ be a diffusion process that gradually adds noise in continuous time $t\in[0,1]$ to ${\mathbf{x}}_{0}$, and let $p_{t}({\mathbf{x}})$ denote the density function of the data at time $t$. Then, this process transforms the real data distribution $p_{0}({\mathbf{x}})$ into a terminal distribution of pure noise $p_{1}({\mathbf{x}})$ from which we can sample. Our goal is to learn the reverse process that allows us to go from noise ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}})$ to a new data sample ${\mathbf{x}}_{0}^{\ast}\sim p_{0}({\mathbf{x}})$.
@@ 2.2 Categorical Features
$\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}\nabla_{{\mathbf{x}}_{t}}\log p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})=\frac{1}{\sigma_{t}^{2}}[\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}[{\mathbf{x}}_{0}]-{\mathbf{x}}_{t}]\ .$ (4)
@@
Likewise, we start the generative process for a categorical feature from an embedding vector ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}})=\mathcal{N}({\bm{0}},\sigma^{2}(1)I_{d})$. We then use score interpolation to iteratively denoise the embedding vector. After the last timestep in the reverse process, that is, $t\approx 0$, we directly infer the generated class from the feature-specific class probabilities, $p({\mathbf{x}}_{0}={\mathbf{e}}_{i}|{\mathbf{x}}_{t},t)$.
This framework can easily be extended to multiple categorical features. All embeddings are trained alongside the model. The model fully takes feature-specific uncertainty at intermediate timesteps into account (for each feature it only has to commit to a category at the final step of generation) which improves the consistency of the generated samples (Dieleman et al., 2022). Therefore, CDTD allows to capture subtle dependencies both within and across data types more accurately.
@@ 3 Method
In short, we combine score matching (Equation 3) with score interpolation (Equation 4) to model the joint distribution of mixed-type data. Next, we discuss the important components of our method. In particular, the combination of the different losses for score interpolation and score matching, initialization and loss weighting concerns, and the adaptive type- or feature-specific noise schedule designs.
@@ 3.1 General Framework
During training, the model predicts the ground-truth value for continuous features and the class-specific probabilities for categorical features. During generation, we concatenate the score estimates, $\hat{s}_{\text{cont}}^{(i)}$ and $\hat{s}_{\text{cat}}^{(j)}$, for all features $i$ and $j$, and pass them to an ODE solver together with $\sigma_{\text{cont},i}$ and $\sigma_{\text{cat},j}$, the noise levels retrieved by transforming linearly spaced timesteps with the learned adaptive noise schedules. Further implementation details are provided in Appendix J.
@@ 3.2 Homogenization of Data Types
Let $\mathcal{L}_{\text{MSE}}(x_{\text{cont}}^{(i)},t)$ denote the time-weighted MSE (i.e., score matching in Equation 3) loss of the $i$-th continuous feature at a single timestep $t$, and $\mathcal{L}_{\text{CE}}(x_{\text{cat}}^{(j)},t)$ the CE (i.e. score interpolation) loss of the $j$-th categorical feature. Naturally, the two losses are defined on different scales. This leads to an unintended importance weighting of features in the generative process. We assume that an unconditional model should a priori, i.e., without having any information, be indifferent between all features. This reflects the state of the model at the terminal timestep $t=1$ in the diffusion process.
@@
The initial equal importance across all timesteps will naturally change over the course of training. We employ an normalization scheme for the average diffusion loss (Karras et al., 2023; Kingma & Gao, 2023) to allow for changes in relative importance among features but ensure equal importance of all timesteps throughout training. To do so, we learn the time-specific normalization term $Z(t)$ such that $\mathcal{L}(t)/Z(t)\approx 1$. This ensures a consistent gradient signal and can be implemented by training a neural network to predict $\mathcal{L}(t)$ alongside our diffusion model (for details see Appendix C).
@@ 3.3 Noise Schedules
Since the optimal noise schedule of one feature impacts the noise schedules of other features, and the different data types have different sensitivities to additive noise, we introduce feature-specific or type-specific noise schedules. For instance, given the same embedding dimension, more noise is needed to remove the same amount of signal from embeddings of features with fewer classes. Likewise, a delayed noise schedule for one feature might improve sample quality as the model can rely on other correlated features that have been (partially) generated first. We make the noise schedules learnable, and therewith adaptive to avoid the reliance on designs for other data modalities.
@@
We aim to learn a noise schedule $h_{k}:t\mapsto\sigma$ for all $K=K_{\text{cont}}+K_{\text{cat}}$ features. Note that $t\in[0,1]$, and with pre-specified minimum and maximum noise levels, we can scale $\sigma_{k}$ to lie in $[0,1]$ as well, without loss of generality. We will learn the feature-specific loss given the noise level, $F_{k}:\sigma_{k}\mapsto\ell_{k}$, alongside the score model, with $\ell_{k}$ the relevant (not explicitly weighted) training loss for the $k$-th feature. Then, our mapping of interest is $h_{k}=\tilde{F}_{k}^{-1}$, that is, the normalized and inverted function $F_{k}$. This encourages the relation between $t$ and $\ell_{k}$ to be linear.
@@
We use the adaptive noise schedules during both training and generation. We derive importance weights from $f_{\mathrm{d.a.log},k}$ to fit $h_{k}$ to avoid biasing the noise schedule to timesteps that are frequently sampled during training. Type-specific noise schedules refer to learning two functions $F_{1}$ and $F_{2}$ that predict the respective average loss over all features of a data type. Examples of learned noise schedules are given in Appendix K.
@@ 4 Experiments
We use a diverse benchmark set of state-of-the-art generative models for mixed-type tabular data. This includes ARF (Watson et al., 2023), CTGAN (Xu et al., 2019), TVAE (Xu et al., 2019), TabDDPM (Kotelnikov et al., 2023), and CoDi (Lee et al., 2023). Each model follows a different design and/or modeling philosophy. Note that CoDi is an extension of STaSy (Kim et al., 2023, the same group of authors) that has shown to be superior in performance. For scaling reasons, the recently proposed Diffusion Forest (Jolicoeur-Martineau et al., 2024) is not an applicable benchmark. Further details on the respective benchmark models and their implementations are provided in Appendix F and Appendix G. To keep the comparison fair, we use the same architecture for CDTD as TabDDPM, with minor changes to accommodate the different inputs (see Appendix J).
@@
We systematically investigate our model on eleven publicly available datasets. The datasets vary in size, prediction task (regression vs. binary classification), number of continuous and categorical features and their distributions. The number of categories for categorical features varies significantly across datasets (for more details, see Appendix E). We remove observations with missings in the target or any of the continuous features and encode missings in the categorical features as a separate category. All datasets are split in train (60%), validation (20%) and test (20%) partitions, hereinafter denoted $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{valid}}$ and $\mathcal{D}_{\text{test}}$, respectively. For classification tasks, we use stratification with respect to the outcome, we condition the model on the outcome during training and generation, and use the train set proportions for generation. In a last post-processing step, we round the integer-valued continuous features after generation for all models.
@@ 4.1 Evaluation Metrics
We evaluate all generative models on four sample quality criteria after generating synthetic data $\mathcal{D}_{\text{gen}}$ of size $\min(\lvert\mathcal{D}_{\text{train}}\rvert,50\,000)$. All metrics are averaged over five random seeds that affect the generative process.
@@
We follow the conventional train-synthetic-test-real strategy (see, Borisov et al., 2023; Liu et al., 2023; Kotelnikov et al., 2023; Kim et al., 2023; Xu et al., 2019; Watson et al., 2023). Hence, we train a group of models, consisting of a (logistic/ridge) regression, a random forest and a catboost model, on the data-specific prediction task (the corresponding hyperparameter settings are reported in Appendix I). We compare the model-averaged real test performance, $\text{Perf}(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$, to the performance trained on the synthetic data, $\text{Perf}(\mathcal{D}_{\text{gen}},\mathcal{D}_{\text{test}})$. We subsample $\mathcal{D}_{\text{train}}$ in case of more than 50 000 observations to upper-bound the computational load. The results are averaged over ten different model seeds (in addition to the five random seeds that impact the sampling process). For regression tasks, we consider the RMSE and for classification tasks, the macro-averaged F1 and AUC scores. We only report $\lvert\text{Perf}(\mathcal{D}_{\text{gen}},\mathcal{D}_{\text{test}})-\text{Perf}(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})\rvert$ in the main part of this paper. An absolute difference close to zero, that is, synthetic and real data induce the same performance, indicates that the generative model performs well.
@@
For each generative model, we report the accuracy of a catboost model that is trained to distinguish between real and generated (fake) samples (Borisov et al., 2023; Liu et al., 2023). First, we subsample the real data subsets, $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{valid}}$ and $\mathcal{D}_{\text{test}}$, to a maximum of 25 000 data samples to limit evaluation time. Then, we construct $\mathcal{D}_{\text{train}}^{\text{detect}},\mathcal{D}_{\text{valid}}^{\text{detect}}$ and $\mathcal{D}_{\text{test}}^{\text{detect}}$ with equal proportions of real and fake samples. We tune each catboost model on $\mathcal{D}_{\text{valid}}^{\text{detect}}$ and report the accuracy of the best-fitting model on $\mathcal{D}_{\text{test}}^{\text{detect}}$ (see Appendix H for details). A (perfect) detection score of 0.5 indicates the model is unable distinguish fake from real samples.
@@
We evaluate the statistical similarity between real and generated training data based on: (1) the Jensen-Shannon divergence (JSD; Lin, 1991) to quantify the difference in categorical distributions, (2) the Wasserstein distance (WD; Ramdas et al., 2017) to quantify the difference in continuous distributions, and (3) the L2 distance of the pair-wise correlation matrices. We use the Pearson correlation coefficient for the correlations between two continuous features, the Theil uncertainty coefficient for the correlations between two categorical features, and the correlation ratio across types (Kotelnikov et al., 2023; Zhao et al., 2021).
@@
That is, the minimum Euclidean distance of a generated data point to any observation in $\mathcal{D}_{\text{train}}$ (Borisov et al., 2023; Zhao et al., 2021). We one-hot encode categorical features and standardize all features to zero mean and unit variance to ensure each feature contributes equally to the distance. We compute the average distance to closest record (DCR) as a robust estimate. For brevity, we report the absolute difference of the DCR of the synthetic data and the DCR of the real test set. A good DCR value, that indicates realistic and sufficiently private data, should be close to zero.

### v3

@@ 1 Introduction
Score-based generative models (Song et al., 2021), also termed diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020), have shown remarkable potential for the generation of images (Dhariwal & Nichol, 2021; Rombach et al., 2022), videos (Ho et al., 2022), text (Li et al., 2022; Dieleman et al., 2022; Wu et al., 2023), molecules (Hoogeboom et al., 2022), and many other highly complex data structures with continuous features. The framework has since been adapted to categorical data in various ways, including discrete diffusion processes (Austin et al., 2021; Hoogeboom et al., 2021), diffusion in continuous embedding space (Dieleman et al., 2022; Li et al., 2022; Regol & Coates, 2023; Strudel et al., 2022), and others (Campbell et al., 2022; Meng et al., 2022; Sun et al., 2023). Diffusion models which include both, continuous and categorical features alike, build directly on advances from the image domain (Kim et al., 2023; Kotelnikov et al., 2023; Lee et al., 2023; Jolicoeur-Martineau et al., 2024) and thus, are not designed to deal with challenges specific to mixed-type data: The different diffusion processes and their losses are neither aligned nor balanced across data types, and do not scale to larger datasets and/or features with a greater number of categories. Models that naively combine different losses to integrate distinct generative processes may suffer from implicitly favoring the sample quality of some features or data types over others (Ma et al., 2020).
@@
A crucial component in score-based generative models is the noise schedule (Kingma et al., 2022; Chen et al., 2022; Chen, 2023; Jabri et al., 2022; Wu et al., 2023). Typical noise schedules for image and text data are designed to focus model capacity on the noise levels most important to sample quality (Nichol & Dhariwal, 2021; Karras et al., 2022), while others attempt to learn the optimal noise schedule (Dieleman et al., 2022; Kingma et al., 2022). For mixed-type tabular data, existing approaches often combine distinct diffusion processes for the continuous and discrete features to derive a joint model (Kotelnikov et al., 2023; Lee et al., 2023). However, noise schedules are not directly transferable from one data modality to another and therefore, using specifications from image or text domain models is not optimal: First, the inherently different diffusion processes make it difficult to balance the noise schedules across features and feature types, and negatively affect the allocation of model capacity across timesteps. Further, and most importantly, the domain, nature and marginal distribution can vary significantly across features (Xu et al., 2019). For instance, any two continuous features may be subject to different levels of discretization or different bounds, even after applying common data pre-processing techniques; and any two categorical features may differ in the number of categories, or the degree of imbalance. The high heterogeneity and lack of balancing warrants a rethinking of fundamental parts of the diffusion framework, including the noise schedule and the effective combination of diffusion processes for different data types.
@@
We counteract the high feature heterogeneity inherent to data of mixed-type with distinct feature or type-specific adaptive noise schedules. The learnable noise schedules allow the model to directly take feature or type heterogeneity into account during both training and generation, and thus avoid the reliance on image or text-specific noise schedule designs. Moreover, we propose a diffusion-specific loss normalization and initialization scheme to homogenize different data types and their losses effectively. Our improvements ensure a better allocation of the model’s capacity across features, feature types and timesteps, and yield high quality samples of tabular data. CDTD outperforms state-of-the-art baseline models across a diverse set of sample quality metrics as well as computation time for data sets with an arbitrary number of categories and data points. Our experiments show that CDTD captures feature correlations exceptionally well, and that explicitly allowing for data-type heterogeneity in the noise schedules benefits sample quality.
@@
  • We propose a joint continuous diffusion model for both continuous and categorical features such that all noise distributions are Gaussian.
  • We balance model capacity across continuous and categorical features with a novel and effective loss calibration, an adjusted score model initialization and type or feature-specific noise schedules.
  • We extend the idea of timewarping and propose a functional form to efficiently learn adaptive noise schedules, and to allow for exact evaluation and easy incorporation of prior information on the relative importance of noise levels.
@@
  • We boost the quality of the generated samples with adaptive, feature or type-specific noise schedules.
  • Our CDTD model allows the first-ever use of advanced techniques, like classifier-free guidance, for mixed-type tabular data directly in data space.
Figure 1: CDTD framework. Adaptive noise schedules are trained to fit the (possibly aggregated) MSE and CE losses and transform the uniform timestep $t$ to a potentially feature-specific noise level to diffuse (“noisify”) the scalar values (for continuous features) or the embeddings (for categorical features). Associated sampling processes are highlighted in orange. The approximated score functions are concatenated and passed to an ODE solver for sample generation.
@@ 2.1 Continuous Features
We denote $x_{\text{cont}}^{(i)}\in{\mathbb{R}}$ as the $i$-th continuous feature and ${\mathbf{x}}_{0}\equiv{\mathbf{x}}_{\text{cont}}\in{\mathbb{R}}^{K_{\text{cont}}}$ as the stacked feature vector. Further, let $\{{\mathbf{x}}_{t}\}_{t=0}^{t=1}$ be a diffusion process that gradually adds noise in continuous time $t\in[0,1]$ to ${\mathbf{x}}_{0}$, and let $p_{t}({\mathbf{x}})$ denote the density function of the data at time $t$. Then, this process transforms the real data distribution $p_{0}({\mathbf{x}})$ into a terminal distribution of pure noise $p_{1}({\mathbf{x}})$ from which we can sample. Our goal is to learn the reverse process that allows us to go from noise ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}})$ to a new data sample ${\mathbf{x}}_{0}^{\ast}\sim p_{0}({\mathbf{x}})$.
@@ 2.2 Categorical Features
$\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}\nabla_{{\mathbf{x}}_{t}}\log p_{0t}({\mathbf{x}}_{t}|{\mathbf{x}}_{0})=\frac{1}{\sigma^{2}(t)}[\mathbb{E}_{p({\mathbf{x}}_{0}|{\mathbf{x}}_{t},t)}[{\mathbf{x}}_{0}]-{\mathbf{x}}_{t}]\ .$ (4)
@@
We start generating a categorical feature from an embedding vector ${\mathbf{x}}_{1}\sim p_{1}({\mathbf{x}})=\mathcal{N}({\bm{0}},\sigma^{2}(1)I_{d})$. We then use score interpolation (Equation 4) to iteratively denoise this embedding vector. After the final reverse step, i.e., when  $t\approx 0$, we can directly infer the generated class from the feature-specific class probabilities, $p({\mathbf{x}}_{0}={\mathbf{e}}_{i}|{\mathbf{x}}_{t},t)$.
This framework can easily be extended to multiple categorical features. All embeddings are trained alongside the model. The model fully takes feature-specific uncertainty at intermediate timesteps into account (as for each feature it only has to commit to a category at the final step of generation) which improves the consistency of the generated samples (Dieleman et al., 2022). Therefore, CDTD allows to capture subtle dependencies both within and across data types more accurately.
@@ 3 Method
In short, we combine score matching (Equation 3) with score interpolation (Equation 4) to model the joint distribution of mixed-type data. Next, we discuss the important components of our method. In particular, the combination of the different losses for score matching and score interpolation, initialization and loss weighting concerns, and the adaptive type- or feature-specific noise schedule designs.
@@ 3.1 General Framework
During training, the model predicts the ground-truth value for continuous features and the class-specific probabilities for categorical features. During generation, we concatenate the score estimates, $\hat{s}_{\text{cont}}^{(i)}$ and $\hat{s}_{\text{cat}}^{(j)}$, for all features $i$ and $j$, and pass them to an ODE solver together with $\sigma_{\text{cont},i}$ and $\sigma_{\text{cat},j}$, the noise levels retrieved by transforming linearly spaced timesteps with the learned adaptive noise schedules. Further details on the implementation and sampling are provided in Appendix J and Appendix K, respectively.
@@ 3.2 Homogenization of Data Types
Let $\mathcal{L}_{\text{MSE}}(x_{\text{cont}}^{(i)},t)$ denote the time-weighted MSE (i.e., score matching) loss of the $i$-th continuous feature at a single timestep $t$, and $\mathcal{L}_{\text{CE}}(x_{\text{cat}}^{(j)},t)$ the CE (i.e., score interpolation) loss of the $j$-th categorical feature. Naturally, the two losses are defined on different scales. This leads to an unintended importance weighting of features in the generative process (Ma et al., 2020). We assume that an unconditional model should a priori, i.e., without having any information, be indifferent between all features. This reflects the state of the model at the terminal timestep $t=1$ in the diffusion process.
@@
The initial equal importance across all timesteps will naturally change over the course of training. We employ a normalization scheme for the average diffusion loss (Karras et al., 2023; Kingma & Gao, 2023) to allow for changes in relative importance among features but ensure equal importance of all timesteps throughout training. To do so, we learn the time-specific normalization term $Z(t)$ such that $\mathcal{L}(t)/Z(t)\approx 1$. This ensures a consistent gradient signal and can be implemented by training a neural network to predict $\mathcal{L}(t)$ alongside our diffusion model (for details see Appendix C).
@@ 3.3 Noise Schedules
Since the optimal noise schedule of one feature impacts the noise schedules of other features, and different data types have different sensitivities to additive noise, we introduce feature-specific or type-specific noise schedules. For instance, given the same embedding dimension, more noise is needed to remove the same amount of signal from embeddings of features with fewer classes. Likewise, a delayed noise schedule for one feature might improve sample quality as the model can rely on other correlated features that have been (partially) generated first. We make the noise schedules learnable, and therewith adaptive to avoid the reliance on designs for other data modalities.
@@
Based on Dieleman et al. (2022), we aim to learn a noise schedule $h_{k}:t\mapsto\sigma$ for all $K=K_{\text{cont}}+K_{\text{cat}}$ features. Note that $t\in[0,1]$, and with pre-specified minimum and maximum noise levels, we can scale $\sigma_{k}$ to lie in $[0,1]$ as well, without loss of generality. We will learn the feature-specific loss given the noise level, $F_{k}:\sigma_{k}\mapsto\ell_{k}$, alongside the score model, with $\ell_{k}$ the relevant (not explicitly weighted) training loss for the $k$-th feature. Then, our mapping of interest is $h_{k}=\tilde{F}_{k}^{-1}$, that is, the normalized and inverted function $F_{k}$. This encourages the relation between $t$ and $\ell_{k}$ to be linear.
@@
We use the adaptive noise schedules during both training and generation. We derive importance weights from $f_{\mathrm{d.a.log},k}$ to fit $h_{k}$ to avoid biasing the noise schedule to timesteps that are frequently sampled during training. Type-specific noise schedules refer to learning two functions $F_{1}$ and $F_{2}$ that predict the respective average loss over all features of a data type. Examples of learned noise schedules are given in Appendix O.
@@ 4 Experiments
We use a diverse benchmark set of state-of-the-art generative models for mixed-type tabular data. This includes SMOTE (Chawla et al., 2002), ARF (Watson et al., 2023), CTGAN (Xu et al., 2019), TVAE (Xu et al., 2019), TabDDPM (Kotelnikov et al., 2023), CoDi (Lee et al., 2023), TabSyn (Zhang et al., 2024). Each model follows a different design and/or modeling philosophy. Note that CoDi is an extension of STaSy (Kim et al., 2023, the same group of authors) that has shown to be superior in performance. For scaling reasons, ForestDiffusion (Jolicoeur-Martineau et al., 2024) is not an applicable benchmark. [^1] Jolicoeur-Martineau et al. (2024) report in the appendix that they used 10-20 CPUs with 64-256 GB of memory for datasets with a median number of 540 observations. With the suggested hyperparameters (for improved efficiency) and 64 CPUs, the model took approx. 500 min of training on the relatively small nmes data. Note that the model estimates $KT$ separate models, with $K$ being the number of features and $T$ the noise levels. Therefore, we consider ForestDiffusion to be prohibitively expensive for higher-dimensional data generation. Further details on the respective benchmark models and their implementations are provided in Appendix F and Appendix G. We provide an in-depth comparison of CDTD to the diffusion-based baselines in Appendix N. To keep the comparison fair, we use the same architecture for CDTD as TabDDPM (the latter has also been adopted by TabSyn), with minor changes to accommodate the different inputs (see Appendix J).
@@
We systematically investigate our model on eleven publicly available datasets. The datasets vary in size, prediction task (regression vs. binary classification [^2] For ease of presentation, we only analyze binary targets. However, CDTD trivially extends to targets with multiple classes.), number of continuous and categorical features and their distributions. The number of categories for categorical features varies significantly across datasets (for more details, see Appendix E). We remove observations with missings in the target or any of the continuous features and encode missings in the categorical features as a separate category. All datasets are split in train (60%), validation (20%) and test (20%) partitions, hereinafter denoted $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{valid}}$ and $\mathcal{D}_{\text{test}}$, respectively. For classification tasks, we use stratification with respect to the outcome, we condition the model on the outcome during training and generation, and use the train set proportions for generation. In a last post-processing step, we round the integer-valued continuous features after generation for all models.
@@ 4.1 Evaluation Metrics
In our experiments, we follow conventions from previous papers and use four sample quality criteria, which we assess using a comprehensive set of measures. All metrics are averaged over five random seeds that affect the generative process, which samples synthetic data $\mathcal{D}_{\text{gen}}$ of size $\min(\lvert\mathcal{D}_{\text{train}}\rvert,50\,000)$.
@@
We follow the conventional train-synthetic-test-real strategy (see, Borisov et al., 2023; Liu et al., 2023; Kotelnikov et al., 2023; Kim et al., 2023; Xu et al., 2019; Watson et al., 2023). Hence, we train a group of models, consisting of a (logistic/ridge) regression, a random forest and a catboost model, on the data-specific prediction task (the corresponding hyperparameter settings are reported in Appendix I). We compare the model-averaged real test performance, $\text{Perf}(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$, to the performance when trained on the synthetic data, $\text{Perf}(\mathcal{D}_{\text{gen}},\mathcal{D}_{\text{test}})$. We subsample $\mathcal{D}_{\text{train}}$ in case of more than 50 000 observations to upper-bound the computational load. The results are averaged over ten different model seeds (in addition to the five random seeds that impact the sampling process). For regression tasks, we consider the RMSE and for classification tasks, the macro-averaged F1 and AUC scores. We only report $\lvert\text{Perf}(\mathcal{D}_{\text{gen}},\mathcal{D}_{\text{test}})-\text{Perf}(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})\rvert$ in the main part of this paper. An absolute difference close to zero, that is, synthetic and real data induce the same performance, indicates that the generative model performs well.
@@
For each generative model, we report the accuracy of a catboost model that is trained to distinguish between real and generated (fake) samples (Borisov et al., 2023; Liu et al., 2023; Zhang et al., 2024). First, we subsample the real data subsets, $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{valid}}$ and $\mathcal{D}_{\text{test}}$, to a maximum of 25 000 data samples to limit evaluation time. Then, we construct $\mathcal{D}_{\text{train}}^{\text{detect}},\mathcal{D}_{\text{valid}}^{\text{detect}}$ and $\mathcal{D}_{\text{test}}^{\text{detect}}$ with equal proportions of real and fake samples. We tune each catboost model on $\mathcal{D}_{\text{valid}}^{\text{detect}}$ and report the accuracy of the best-fitting model on $\mathcal{D}_{\text{test}}^{\text{detect}}$ (see Appendix H for details). A (perfect) detection score of 0.5 indicates the model is unable distinguish fake from real samples.
@@
We aim to assess the statistical similarity between real and generated data at both the feature and sample levels. We largely follow Zhao et al. (2021) and compare: (1) the Jensen-Shannon divergence (JSD; Lin, 1991) to quantify the difference in categorical distributions, (2) the Wasserstein distance (WD; Ramdas et al., 2017) to quantify the difference in continuous distributions, and (3) the L2 distance between pair-wise correlation matrices. We use the Pearson correlation coefficient for two continuous features, the Theil uncertainty coefficient for two categorical features, and the correlation ratio for mixed types. Similar metrics for the evaluation of statistical similarity have been used by Zhang et al. (2024).
@@
That is, the minimum Euclidean distance of a generated data point to any observation in $\mathcal{D}_{\text{train}}$ (Borisov et al., 2023; Zhao et al., 2021). We one-hot encode categorical features and standardize all features to zero mean and unit variance to ensure each feature contributes equally to the distance. We compute the average distance to closest record (DCR) as a robust estimate. For brevity, we report the absolute difference of the DCR of the synthetic data and the DCR of the real test set. A good DCR value, indicating both realistic and sufficiently private data, should be close to zero.